# iTBLS: A Dataset of Interactive Conversations Over Tabular Information

### v1

@@ 5 Results
We introduce a modular approach where we use an LLM to identify the nature of the interaction followed by another LLM to execute the operation described in the user utterance. A high-level overview is depicted in Figure 3. We use the version of Llama 2 Touvron et al. (2023) already pre-trained and instruction-aligned using supervised fine-tuning and reinforcement learning with human feedback for dialogue use cases (Llama-2-[7/13/70]B-chat-hf) [^3] https://huggingface.co/meta-llama, hereafter referred to as Llama-2-[7/13/70]B.
Figure 3: A block diagram of our approach to iTBLS
5.1 Interaction Identification
For interaction identification we benchmark both zero-shot and fine-tuned approaches using GPT-{3.5,4} and Llama 2-{7B,13B,70B}. We use classification accuracy as the metric. For zero-shot experiments, we instruct the model to wrap the identified task in special tags, e.g. “[TASK] modify [/TASK]". This enables automated extraction of the answers in the zero-shot setting. Any response that fails to follow the requested format is considered incorrect. We aggregate all data across tasks into pairs of requests $i$ and tasks $k$ $(r_{i,k},t_{k})$ and maximize the objective:
$p(t_{k}|r_{i,k})$ (1)
Results for interaction identification are provided in Table 2 where PEFT using Llama 2 achieves near-perfect performance on the task.
Model Meth. Split Accuracy
Llama-2-7b ZS dev 0.531
test 0.516
Llama-2-13b ZS dev 0.163
test 0.150
Llama-2-70b ZS dev 0.501
test 0.483
GPT-3.5 turbo ZS dev 0.431
test 0.416
GPT-4 ZS dev 0.658
test 0.658
Llama-2-7b PEFT dev 0.997
test 0.997
Table 2: Interaction ID: Zero-shot (ZS) vs PEFT
5.2 Tabular Interpretation
Table 3 contains results for the zero-shot approach using the different sizes of Llama 2. For each model, we report the exact match score that identifies whether the entire answer string is present in a continuous form in the model’s response. Using larger models results in better exact match scores, with the answer phrase present in almost 30% of the responses from the 70B variant of Llama 2. We also report results using PEFT. In this situation, Llama 2 is fine-tuned to maximize the probability of the correct answer over the set of all possible answers $\mathcal{A}$ given a table and a user request:
$\operatorname*{argmax}_{\mathcal{A}}p(\text{answer}|\text{table},\text{request})$ (2)
Hyperparameters for fine-tuning are detailed in Appendix A.2. As expected, fine-tuning using QLoRA performs better than zero-shot prompting, achieving up to 2x the exact match score when compared to the zero-shot approaches (Table 3). Our best-performing method is a multi-step approach combined with parameter-efficient fine-tuning. In this method, we identify the cell referred to by the user utterance in the concept of the table (row and column number) and then generate a response based on this cell. Performing this operation ensures that the output response is necessarily a cell of the table as opposed to free-form generation with no constraint on the emitted LLM response being a cell of the input table. An example of the codified representation is available in Figure 2.
Model Method Split Exact
Match
Llama-2-7b ZS dev 0.176
test 0.199
Llama-2-13b ZS dev 0.170
test 0.213
Llama-2-70b ZS dev 0.233
test 0.291
Llama-2-7b PEFT dev 0.517
test 0.560
Llama-2-7b MS+PEFT dev 0.667
test 0.631
Table 3: Results: iTBLS-interpretation
5.3 Tabular Generation
First, we experiment with zero-shot prompting with Llama 2 to incorporate the row or column information into the table in a single step. During evaluation, we assume tabular information is order invariant, provided the contents align accurately with the correct header cells. To assess the effectiveness of our approach, we report the F1 score, that is, the harmonic mean of precision and recall computed between the cells of the predicted row or column and the ground truth. As evidenced by the results in Table 5, this single-step zero-shot prompting results in tables that are malformed, that is, the generated tables are inconsistent and incorrectly formatted, or do not match the dimensions of the expected ground truth to be generated. The malformed rate reduces as we experiment with larger models, going down from 88% to $\approx$ 26% with a marginal improvement in F1 scores.
To address this issue, we experiment with a multi-step approach. Prior work has shown the effectiveness of reformulating the table-generation task as two-stage machine reading comprehension Sundar et al. (2024) to achieve tables that are always well-formed. The first stage identifies whether the input utterance asks to generate a row or column using Llama 2 in a zero-shot configuration. Then, based on whether a row or column is to be generated, questions are synthesized using the header rows of the corresponding column or row. The synthesized questions are then answered using Llama 2 in a zero-shot configuration to yield the final output. Since we constrain the generated content to conform to the existing tabular structure, none of the tables contain formatting errors, resulting in a malformed rate of 0%.
To further improve results, we experiment with PEFT using QLoRA. The fine-tuning process is performed to ensure the LLMs emit cells of a table autoregressively by optimizing the objective in Equation 3. The output is parsed to reformat the CSV-serialization back into the original two-dimensional representation.
$\operatorname*{argmax}_{h}p(\text{cell}_{n}|\{\text{cell}_{ij}\},\text{table},\text{request})$ (3)
This results in a 2x improvement in F1 scores over the multi-step approach (Table 5), but at the cost of increased malformed rates. To alleviate this issue, we combine the multi-step and fine-tuning approaches. First, Llama 2 is used in a zero-shot configuration to determine whether a row or column is to be added to the table. Next, the question-answering reformulation is used to synthesize question-answer pairs. Llama 2 is then fine-tuned using QLoRA to answer these questions. This approach marries the best of both worlds, resulting in a further 2x improvement in F1 scores over the end-to-end fine-tuning while bringing the malformed rates to 0%.
5.4 Tabular Modification
Model Method Split Acc. Malf. Rate(%)
Llama-2-7b ZS dev 0.006 64.7
test 0.000 14.0
Llama-2-13b ZS dev 0.014 44.1
test 0.002 12.9
Llama-2-70b ZS dev 0.037 15.6
test 0.031 8.1
Llama-2-7b PEFT dev 0.139 2.5
test 0.140 3.1
Llama-2-7b MS+ PEFT dev 0.274 0
test 0.328 0
Table 4: Results: iTBLS-modification
For the modification task, we experiment with a fine-tuning approach with Llama-2-7b, as well as zero-shot approaches with three llama-2 models (7b, 13b, and 70b). For each experiment, we compute table modification accuracy. An output table is accurate when all row and column headers match the ground truth, as well as all cell values within the table. This metric is computed by converting the textual form of the output table into a DataFrame using the pandas library in python and then comparing it to ground truth. If the conversion to DataFrame fails, the table is considered malformed.
The accuracy and malformed rate are reported in Table 4. Zero-shot approaches with all three models achieve low (<4%) accuracy on the modification task, including a 0% accuracy from the smallest model (7B). This is not unexpected due to the need to identify cells from relative references in natural language text form and perform the correct swap, combined with the strict constraint of generating well-formed tables. All this must be performed while maintaining the remainder of the table’s values. This is further exacerbated by the presence of malformed table outputs, which make up as much as 15.6% of output tables in some experiments. Table 4 further demonstrates the strength of PEFT, where the standard approach boosts accuracy by 10.9% on the test split while using 1/10 the parameters. However, malformed tables remain a problem even with standard PEFT. This is addressed with a multi-step approach. First, we generate an encoding of the requested modification. This is then followed by a program synthesis to perform the swap on the table and return the resulting modified table. An example of the encoding to swap contents in the 3rd and 4th columns of the first row is: modify<SEP>1<SEP>4<SEP>1<SEP>3. The corresponding program synthesis is executed in pandas using df.iloc [^4] https://pandas.pydata.org/pandas-docs/stable/reference/api/pandas.DataFrame.iloc.html.
With our multi-step approach, we are able to achieve a 0% malformed rate (this is guaranteed by the method), and an 18.8% absolute increase in accuracy on the test split, thus demonstrating the value of the multi-step approach for the task of tabular modification.
Model Meth. Split F1 Malf. Rate(%)
Llama-2-7b ZS dev 0.000 87.9
test 0.000 87.9
Llama-2-13b ZS dev 0.003 57.8
test 0.005 62.1
Llama-2-70b ZS dev 0.063 29.5
test 0.078 22.6
Llama-2-7b MS dev 0.163 0.0
test 0.128 0.0
Llama-2-13b MS dev 0.094 0.0
test 0.073 0.0
Llama-2-70b MS dev 0.177 0.0
test 0.129 0.0
Llama-2-7b PEFT dev 0.309 12.0
test 0.315 4.9
Llama-2-7b MS+ PEFT dev 0.682 0.0
test 0.699 0.0
Table 5: Comparison of approaches for iTBLS-generation, ZS=Zero-shot, MS=Multi-step, PEFT=Parameter-efficient Fine-tuning
@@ 6 Conclusion
This paper introduces Interactive Tables (iTBLS), a dataset of interactive conversations situated over tables in scientific papers. iTBLS consists of multi-turn conversations with three modes of interaction - interpretation, modification, and generation. iTBLS extends prior work by introducing new interactions such as comparison of table contents, mathematical and positional reasoning, manipulation using natural language commands, and generation of content from new evidence. Additionally, a set of baseline approaches for a variety of computing situations addressing iTBLS is presented. The baseline approaches utilize zero-shot prompting and parameter-efficient fine-tuning. Furthermore, a novel multi-step task-specific approach is introduced. Leveraging the multi-step process in conjunction with parameter-efficient fine-tuning yields the best performance on iTBLS, outperforming standard parameter-efficient fine-tuning by up to 15% on interpretation, 18% on modification, and 38% on generation.
7 Limitations
@@ Appendix A Appendix
A.1 Dataset Examples
Example
1 What is the 2nd cell value for row 4?
2 Tell me the final value in the column labeled k
3 What is the value of the cell to the left of the cell in the bottom right of the table.
Table 6: Example interactions in the Interpretation task of iTBLS
Example
1 The rows 1 and 4 in the Column “Citation" were accidentally switched. Please rectify the positions of these values so they are where they need to be.
2 Swap the contents of the second and last cell under repetitions.
3 Two values in the MCBLp column were put in the reverse spots. I need the values for the FM and PCC rows flipped.
Table 7: Example interactions in the Modification task of iTBLS
Example
1 The row 3 of the table shows the values for Peak as 4, X coordinate as 0.100, Y coordinate as -0.150, A as 0.5, standard deviation ($\sigma$) as 0.02, and Local lnZ as -7.824.
2 The column “Method 2 (with sub-clustering)" contains the ‘Nlike’ values in different rows: 27,658 in the second row, 69,094 in the third row, 579,208 in the fourth row, and 43,093,230 in the fifth row, while the remaining rows from six to nine contain no data (NaN).
3 The column R contains eight numerical values in increasing order: 3.34, 3.40, 3.66, 5.06, 6.02, 6.61, 4.05, and 4.11.
Table 8: Example interactions in the Generation task of iTBLS
A.2 Hyperparameters
Parameter Efficient Fine-tuning
Parameter Value
Rank 64
$\alpha$ 16
Dropout 0.1
Table 9: Hyperparameters for Parameter-efficient fine-tuning using QLoRA
Additional hyperparameters:
Parameter Value
Learning Rate 2e-4
Batch size 4
Warmup Schedule Constant
Warmup Ratio 0.03
Epochs 5
Optimizer paged_adamw_32bit [^5] https://huggingface.co/docs/bitsandbytes/main/en/reference/optim/adamw
Compute 8 Nvidia A40 GPUs
Table 10: Additional hyperparameters for fine-tuning experiments
A.3 Mechanical Turk Interface
Figure 4: Amazon Mechanical Turk Interface to collect iTBLS interpretation
Figure 5: Amazon Mechanical Turk Interface to collect iTBLS modification

### v2

@@ 5 Results
5.1 Experimental Setup
For our experiments, we utilize Gemma models Team et al. (2024). We fine-tune the instruction-tuned base model gemma-2-9b-it using LoRA Hu et al. (2022). Hyperparameters for our training setup as well as LoRA parameters are shown in Appendix A.
5.2 Datasets
In addition to the iTBLS dataset, we also evaluate our method on five datasets to summarize textual paragraphs to tables Wu et al. (2022). While iTBLS is a table-to-table or table-to-text task, the datasets proposed by Wu et al. (2022) address the dual problem of text-to-table. The datasets consist of textual paragraphs containing some information that is to be converted into a tabular format by determining both the appropriate header cells and the content that the table is filled with.
Wu et al. (2022) present datasets for the text-to-table task by inverting datasets created for the dual problem of generating textual descriptions from tables. Each dataset consists of textual paragraphs paired with tabular information summarizing content in the text. Dataset statistics are available in Appendix A.3. Each dataset is described below.
E2E Novikova et al. (2017) concerns restaurant descriptions, requiring summarization of information into tables with descriptors like restaurant name, customer rating, and location. WikiTableText (WTT) Bao et al. (2018), sourced from Wikipedia, consists of natural language descriptions generated from tabular data across various topics. WikiBio Lebret et al. (2016) comprises introductions of individuals from Wikipedia alongside tabular summaries extracted from the same page’s information box. In contrast to E2E, the table headers in the WikiTableText and WikiBio datasets vary widely across data samples.
Example textual paragraphs and associated tables from each dataset are presented in A.4.
5.3 Metrics
Exact-Match (EM): On the iTBLS dataset, we report exact-match, that is, whether or not the generated table matches the ground-truth table exactly.
BERTScore: On the E2E, WTT, WikiBio and RotoWire datasets, we report BERTScore Zhang et al. (2020) in addition to EM to be consistent with prior work. BERTScore is a measure of semantic similarity which computes the similarity of embeddings in a latent space obtained using an encoder language model.
Consistent with prior work, all our evaluations are order-invariant. That is, credit is given as long as the generated cells are indexed by the correct row and column headers, even if the headers themselves are in different positions between the model-generated response and the ground-truth.
5.4 iTBLS
Results on the iTBLS dataset using a vanilla sequence-to-sequence approach and the question-answering-based method are presented in Table 2. As observed in the results, the generate task is the hardest, with performance slightly lower on the generate task when compared to interpret and modify. This is a result of the fact that the exact-match metric only provides credit when all cells are correct (necessitating that all cells in the output are identical to the ground truth) and does not provide partial credit for getting some of the cells right, and the fact that the generate task requires getting more cells right in comparison to the other tasks.
Split Approach Exact-Match
Interpret Seq2seq 88.29
iTBLS as QA 90.98
Modify Seq2seq 74.65
iTBLS as QA 89.58
Generate Seq2seq 48.94
iTBLS as QA 73.32
Table 2: Comparison between the question-answering reformulation and a vanilla sequence-to-sequence modeling approach on the iTBLS dataset
5.5 Text-to-table
The results on the text-to-table datasets proposed by Wu et al. (2022) are available in Table 3. Our method performs on par with or better than the prior state-of-the-art method in terms of BERTScore and is competitive with prior work in terms of Exact-Match. The exact-match score does not reflect true performance on the WikiBio dataset since synonyms are penalized under this framework. A deep-dive into the results is presented in Section 5.6.
Dataset Approach EM BS
WTT Wu et al. (2022) 62.71 80.74
Ours 75.96 95.52
Wikibio Wu et al. (2022) 69.71 76.56
Ours 66.65 92.60
E2E Wu et al. (2022) 97.94 98.57
Ours 97.64 99.35
Table 3: Comparison between our method and prior work on the text to table task in terms of Exact-Match and BERTScore
5.6 Analysis of Errors
An analysis of the difference in performance between the prior state of the art and our approach is presented in Table 6. As observed, the dataset is inconsistent in the description of individuals, with no consistent pattern when middle and last names are present. Furthermore, the use of quantifying information in the header as opposed to the table cell results in no credit using the exact-match metric, though the information contained is exactly the same between the prediction and the ground-truth. Finally, the datasets often contain textual examples with multiple possible tabular summarizations, all of which are equally valid, further complicating evaluation. In the third example in Table 6, the model correctly generates the ‘Occupation’ as a table header while the ground truth contains an erroneous sample, using the phrase ‘Known for’ instead of ‘Known as’.
Examples of errors in the iTBLS dataset are provided in Tables 4, 5, and LABEL:tab:difference_answers_itbls_generate. On the interpret task, the model incorrectly understand the user request, and produces the cell immediately to the right instead of three columns over. On the modify task (Table 5), the model incorrectly understands the references and swaps index (2,3) with (3,2) instead of swapping indices (2,2) and (3,3). On the generate task (Table LABEL:tab:difference_answers_itbls_generate), the model incorrectly places a tuple and hallucinates a value instead of performing the requested action.
Text: What is the value of the cell in row 1 that is three cells to the right of the cell with a value of 12%?
Input Table: row ID $\sigma\mu[I0]$ $\mu[\tau s]$ $\sigma[\tau s]$ $\sigma\mu[\tau s]$ 0 13% 912.5 $\mu s$ 91.9 $\mu s$ 10.1% 1 12% 18335.7 $\mu s$ 90.7 $\mu s$ 10.0% 2 12% 903.1 $\mu s$ 1832.7 $\mu s$ 10.0% Ground Truth: 10.0% Prediction: 18335.7 $\mu s$
Table 4: Example error for iTBLS interpret task. Table source: https://arxiv.org/pdf/1411.5458
Text: Swap row 1 in the second column with row 2 in the third column
Input Table: row ID col 1 col 2 col 3 0 X O X 1 NaN O O 2 O X X Ground Truth: row ID col 1 col 2 col 3 0 X O X 1 NaN X O 2 O X O Prediction: row ID col 1 col 2 col 3 0 X O X 1 NaN O X 2 O O X
Table 5: Example error for iTBLS modify task. Table source: https://arxiv.org/pdf/1411.4023
1. Text: Walter Clarence Henderson (28 February 1891 – 20 September 1968) was a progressive conservative party member of the Canadian house of commons. He was born in Carberry, Manitoba and became a farmer by career. He was elected at the Cariboo riding in the 1958 general election, defeating social credit incumbent Bert Leboe.
Generated Table:
Predicted Header Prediction - iTBLS Ground Truth Header Ground Truth
Name Walter Henderson Name Walter Clarence Henderson
Profession Farmer Profession Farmer
Party Progressive Conservative Party Progressive Conservative
2. Text: The production of Tautona mine is 235,000 ounces in 2013.
Generated Table:
Predicted Header Prediction - iTBLS Ground Truth Header Ground Truth
Title Tautona mine Title Tautona mine
Subtitle Production Subtitle Production
Year 2013 Year 2013
Production (ounces) 235,000 Production 235,00 ounces
3. Text: Elango Nagarajah, also known as “Thaimann Elango", is a Tamil film actor, director, producer and lyricist in the Tamil film industry. He began his career in his early ages as a producer for the Tamil film Anbudan, starred Arun Vijay, Meena, Rambha (actress) in the main was released in the year 2000.
Generated Table:
Predicted Header Prediction - iTBLS Ground Truth Header Ground Truth
Name Elango Nagarajah Name Elango
Occupation actor, director, producer, lyricist Known for Thaimann
Table 6: Difference between the tables generated by the Zero Shot (ZS) and Fine-Tuned (FT) approaches with respect to the Ground Truth on the WikiBio and WikiTableText datasets with additions and deletions represented using red and green.
@@ 6 Conclusion
This paper introduces Interactive Tables (iTBLS), a dataset of interactive conversations addressing three types of tasks – interpretation, modification, and generation. In contrast to prior tabular datasets that are sourced from Wikipedia or financial reports, iTBLS is situated in tabular data obtained from scientific pre-prints on ArXiv. Success on the iTBLS dataset requires understanding both ordinal and cardinal references to cell positions, and understanding implicit references. Additionally, the paper introduces a novel framework that reformulates tabular operations as question-answering. Appropriate questions are created based on the input table and the nature of interaction, and the user request is used as evidence to obtain the answers. The developed approach demonstrates an improvement over a sequence-to-sequence modeling approach on the iTBLS dataset. In addition, the question-answering-based reformulation is evaluated on datasets for the text-to-table task, obtaining up to 13% improvement in terms of exact-match accuracy and 16% improvement in terms of BERTScore compared to the prior state-of-the-art.
Limitations
@@
Acknowledgments
This work was supported by CoCoSys, one of seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA.
@@ A.1 AI Assistance Acknowledgment
A.2 Compute
All fine-tuning and inference was run on Nvidia A40 GPUs with 48GB GDDR6 memory. Fine-tuning took  1-2 hours on 8 GPUs in parallel with pytorch distributed data parallel (DDP).
A.3 Dataset Statistics
Statistics of the text-to-table datasets:
Dataset Train Valid Test
E2E 42.1k 4.7k 4.7k
WikiTableText 10k 1.3k 2.0k
WikiBio 582.7k 72.8k 72.7k
Table 7: Statistics of the E2E, WikiTableText, WikiBio, and RotoWire datasets, number of samples across splits
A.4 Dataset Examples – Text to Table
This section details example textual paragraphs and associated tables from the different datasets.
E2E:
The Eagle is a low rated coffee shop near Burger King and the riverside that is family friendly and is less than £20 for Japanese food.
Name The Eagle
Food Japanese
Price range Less than £20
Customer Rating Low
Area Riverside
Family friendly Yes
Near Burger King